What is the factored expression of x^2+21x-100
Answer by josgarithmetic(39633)

Keep on using what you have been learning. What two numbers, integers, have product -100 and have sum +21?

Pick possibilities using the product 100.
10 and 10
5 and 20
4 and 25
2 and 50
One of the parts of the combination will need (most likely) to be negative and the other positive.

Look at 4 and 25. ONE of them must be positive and THE OTHER must be negative; but which?
-4+25= what?
4+(-25)=what?
(x_______)*(x_______)=x^2+21x-100